How would a spinning disk's kinetic energy change
Answer #1

Rotational K.E = 0.5*I*w^2

I= moment of inertia

and w= angular speed


when I' = 5*I

and w' = w / 5

=> K.E ' = 0.5 *5*I*(w/5)^2

=> K.E ' = 0.2 * (0.5*I*w^2)


=> K.E ' = 0.2 K.E hence option (E) ans...